OpenStudy (anonymous):

Find the identical expression for the following: sin(pi/4 +x) + sin(pi/4-x)

OpenStudy (anonymous):

is it sqrt(2)*cosx

OpenStudy (anonymous):

how do you get that?

OpenStudy (chaise):

I think so - but don't ask me how ^_^

OpenStudy (anonymous):

use sin(a+b)+sin(a-b)=2sina cosb

OpenStudy (anonymous):

sin(a+b) = sina cosb + cosa sinb sin(a-b) = sina cosb - cosa sinb (sina cosb + cosa sinb)+(sina cosb - cosa sinb)

OpenStudy (anonymous):

= 2sina cosb

OpenStudy (anonymous):

thank you